OpenStudy (anonymous):

solve -9y-12(2-3y)=-25-2(y+6)

OpenStudy (anonymous):

Multiply Out the Brackets -9y-24+36y=-25-2y-12 Get all the y values on one side and the constant on the other -9y+36y+2y=-25-12+24 29y=-13 y=-13/29

OpenStudy (anonymous):

thank you so much, I just couldnt get this one

OpenStudy (anonymous):

no problem, glad to have helped

OpenStudy (anonymous):

can anyone help me with this problem

OpenStudy (anonymous):

what problem?

OpenStudy (anonymous):

100-3.4 4 to the second power divided by 2 to the third power -(-5) to the second power

OpenStudy (anonymous):

sorry im not sure how to write it all out , i gave it my best shot

OpenStudy (anonymous):

if u put one of these ^ before the power i will be able to understand it better, if u write the question with the powers

OpenStudy (anonymous):

100-3.4^ divided by 2^ -(-5)2^

OpenStudy (anonymous):

like that? this is my first algebra class and I am totally lost

OpenStudy (anonymous):

It's ok, dont worry i just dont want to read the question wrong and give you a wrong answer!

OpenStudy (anonymous):

thanks, I have to graph charts as well on a different problem and i have no clue how to do it. the teacher is very vague and doesnt offer much help at all

OpenStudy (anonymous):

im trying to think of the best way you can write it if u are unsure. It's fairly difficult to read otherwise

OpenStudy (anonymous):

okay thank you so much

OpenStudy (anonymous):

If i were to say 3 to power of 4/3 was written like this: 3^(4/3) are you able to write your question in a similar way?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, can u write it like this then please?

OpenStudy (anonymous):

i will try

OpenStudy (anonymous):

100-3.4 __________ 4^2 divided by 2^-(-5)2^ the line between the top eauation and bottom one is suppose to be solid